# Real-Time Recurrent Learning using Trace Units in Reinforcement Learning

### v1

@@ 4.1 Ablation Study on Architectural Choices for RTUs and LRUs
In this first experiment, we investigate the impact of several architectural choices on the performance of RTUs and LRUs varying where nonlinearity is applied. We use a simple multi-step prediction task called Trace conditioning [38] inspired by experiments in animal learning. The agent’s objective is to predict a signal—called the US—conditioned on an earlier signal—-the CS. The prediction is formulated as a return, $G_{t}\doteq\sum_{k=0}^{\infty}\gamma^{k}{\text{US}}_{t+k+1}$, where the agent’s goal is to estimate the value function for this return. More details on this environment and experimental settings are in Appendix E. Figure 1 summarizes the results.
@@ 4.2 Learning under resources constraints
In the following experiments, all agents consist of a recurrent layer followed by a linear layer generating the prediction. We measure performance of the agents online by calculating the Mean Square Return Errors (MSRE), which is the mean squared error between the agent’s prediction at time $t$ and $G_{t}$. In all the experiments, we used the Adam optimizer. We first ran each agent with different step sizes for $5$ runs and $2$ million steps. We then averaged the MSRE over the $2$ million steps and selected each agent’s best step size value. Finally, we ran the agents with the best step size value for $10$ runs, which we report here. We also report all agents’ step size sensitivity curves in Appendix E.3.
@@
We provide additional experiments comparing RTUs to two other approaches that use RTRL: online LRUs and a real-valued diagonal RNN in Appendix E.1.
@@ 5.1 Linear RTRL Methods in Incremental and Batch Settings
Integrating Linear RTRL Methods with PPO When performing batch updates, as with PPO, the RTRL gradients used to update the recurrent parameters will be stale, as they were calculated during the interaction with the environment w.r.t old policy and value parameters. One solution to mitigate the gradient staleness is to go through the whole trajectory after each epoch update and re-compute the gradient traces. However, this can be computationally expensive. In Appendix F, Algorithm 1, we provide the pseudocode for integrating RTRL methods with PPO with optional steps for re-running the network to update the RTRL gradient traces, the value targets, and the advantage estimates. We also performed an ablation study to investigate the effect of the gradient staleness in RTRL when combined with PPO, Appendix F. The results from the ablation study show that using a stale gradient results in better performance with RTUs and suggests that the staleness might help PPO maintain the trust region.
@@ 6 Experiments in Memory-Based Control
We evaluated our Linear and Non-linear RTUs against GRU, LRU, and Online LRU. All agents use PPO [39] as the control algorithm, and the difference between the agents is the recurrent component. For all agents, we fixed the number of parameters for the recurrent part to be $\sim 24$k. We tuned the learning rate for all agents in all environments and selected the best learning rate for each agent per environment. We also included a GPT2-transformer baseline. We followed the implementation details in previous work  [31], and used a GPT2 variant with $200$k parameters. We provide the results for GPT2 in Appendix G.
@@
When given the same number of parameters, RTU agents outperform other baselines in all environments in Figure 4. Furthermore, we show in Appendix G that even when increasing the truncation length of both GRU and LRU agents to use significantly longer history, they outperform RTUs in only one task. Of particular note is again that RTUs outperform online LRUs, highlighting again that our simple modifications have a large impact on performance in this online RL setting.
@@ 8 Acknowledgments
Appendix A More Details on Representability with Complex-valued Diagonal Recurrence
@@
A.1 Representability with Complex-valued Diagonals
@@
The above equivalence has only been used without any activation, namely in linear-recurrent units (LRUs) [33]. A natural question is if only the identity $f(\mathbf{x})=\mathbf{x}$ (and linear functions) satisfy this property of commuting with eigenvector matrices. Intuitively, this seems like the only option, as imagining a nonlinearity that commutes is hard. Surprisingly, for the more restricted case of symmetric $\mathbf{W}_{h}$, we can show a slightly more general class of activations can be used, proving an if-and-only-if relationship (see Appendix A.2). However, even for this restricted setting, this generalized class is limited and such activations unlikely to be preferable to a linear recurrence. We see this as a negative result, that suggests this equivalence only holds for the linear setting.
A.2 More on the Equivalence of Non-Linear RTUs and Dense RNNs
@@
A.3 Complex Eigenvalues in Vanilla RNNs
We empirically investigate whether complex eigenvalues appear when training dense RNNs in a simple task. The goal is to show that $\mathbf{W}_{h}$ is not a symmetric matrix, even in the simplest tasks, so having only real-valued diagonals is too restrictive. We used the Three State POMDP[41] for this experiment.
The Three State MDP is depicted in Figure 7. In this task, the agent needs to remember one cue from the previous time-step ago, to make a prediction about the next time-step. The MDP has three states, $s_{1}$, $s_{2}$, and $s_{3}$, and no actions. If the agent is in either $s_{1}$ or $s_{2}$, it transitions to any of the three states with equal probability. However, if the agent is in $s_{3}$, it transitions to the state preceded by $s_{3}$. A sequence of observations would look like $1,3,1,2,2,3,2,\cdots$, and we ask the agent to predict the next observation.
Appendix B Background on BackPropagation Through Time and Real-Time Recurrent Learning
B.1 BackPropagation Through Time
Applying the chain rule, we re-write Eq.13 as:
Once again, we need to consider the two pathways when evaluating $\frac{\partial\mathbf{h}_{t-1}}{\partial\boldsymbol{\psi}}$ in 16. For simplicity, let $\mathbf{J}_{t}\doteq\frac{\partial\mathbf{h}_{t}}{\partial\boldsymbol{\psi}}$, $\mathbf{B}_{t}=\frac{\partial\mathbf{f}(\mathbf{h}_{t-1},\mathbf{x}_{t},\boldsymbol{\psi})}{\partial\boldsymbol{\psi}}$, $\mathbf{C}_{t}=\frac{\partial\mathbf{f}(\mathbf{h}_{t-1},\mathbf{x}_{t},\boldsymbol{\psi})}{\partial\mathbf{h}_{t-1}}$, and re-write 16:
Writing $\frac{\partial\mathbf{h}_{t}}{\partial\boldsymbol{\psi}}$ using the results from 17:
B.1.1 Truncated-BackPropagation Through Time
Williams and Peng (1990) introduced Truncated-BackPropagation Through Time (T-BPTT) which solves the issue of increasing memory and computational complexities of BPTT. In T-BPTT, we specify a truncation length $T$, which controls the number of steps considered when calculating the gradient in 18. We now write the truncated version of 17 which takes into consideration the gradient from the last $T$ steps only:
Using results from 19, we then write the approximated gradient of the loss w.r.t the learnable parameters:
B.2 Real-Time Recurrent Learning
Williams and Zipser (1989) introduced the Real-time Recurrent Learning algorithm (RTRL) as a learning algorithm for continual recurrent learning. RTRL employs the recurrent formulation of the gradient in 16; instead of unrolling $\frac{\partial\mathbf{h}_{t-1}}{\partial\boldsymbol{\psi}}$ further back in time, RTRL saves its calculated value from the previous time step and use it later when needed. It is worth emphasizing that after the agent updates its parameters, the gradient information saved from previous time steps would be stale, i.e., calculated w.r.t old parameters, however, under the assumption of small learning rates, RTRL is known to converge. The gradient formulation of RTRL can be written as:
Appendix C Recurrent Trace Units
C.1 Empirical Analysis for different $r$ and $\theta$ parameterizations:
Since $r$ represents the magnitude of a complex number, then $r\in\mathbb{R}^{+}$. Additionally, it is preferred to have $r\in(0,1]$ to avoid vanishing/exploding gradients. Let $w_{r}$ be a learnable parameter which could directly represent $r$ or represent a function of $r$, we can enforce the constraints on $r$ in several ways:
  1. 1.
  2. 2.
  3. 3.
C.2 Real-Time Recurrent Learning for a Single Layer Linear RTUs
We can derive the following gradients:
where
We then get:
C.3 Single Layer Non-Linear RTUs Formulation:
We then get:
C.4 Complexity Analysis of RTUs
The saved gradient information has the following dimensions:
For the computational complexity, a forward pass according to 2 has a computational complexity of $\mathcal{O}(n+nd)$. Additionally, after doing the forward pass, the learner needs to update the saved gradient information according to equations 23 through 27 which has a computational complexity of $\mathcal{O}(n+nd)$. To summarize, using Real-Time Recurrent Learning with RTUs has linear computational and memory complexities.
C.5 Multi-Layers Recurrent Trace Units
where
C.6 Implementing RTRL within the reverse-mode automatic differentiation
@@ D.2 Biased gradient when only using the real part of the hidden state
Appendix E Trace Conditioning Experiments
@@
E.1 Additional Trace Conditioning Experiments
@@
The results in figure E.1 indicate that these seemingly small differences between the diagonal RNNs can result in significantly different behavior. RTUs outperform online LRUs, with the differences discussed in-depth in Section 3.4. RTUs also outperform the block diagonal RNN. We emphasized using real-valued diagonals implicitly assumes symmetric matrices, but that is for a single real-value. This block diagonal has more representational capacity than the RTU. This result suggests it is beneficial for learning to enforce these constraints on the learnable parameters, that they correspond to the rotational representation of complex numbers.
On the role of RTRL in RTUs: We investigated the importance of using RTRL rather than T-BPTT for RTUs. We found that RTUs with T-BPTT perform similarly to LRUs with T-BPTT. With a sufficiently large truncation, performance eventually matched that of RTRL, but with significantly more computation. See Appendix E.1 for the results.
@@
To highlight the role of RTRL in RTUs, we evaluated the performance of both linear and non-linear RTUs with T-BPTT. In this experiment, we all agents use the same number of parameters; the only difference is whether they use RTRL or BPTT.
E.2 Trace Conditioning Experiments Details
@@
E.3 Learning Rate Sensitivity
@@
Figure 16: RTUs used in the animal learning benchmark.
Appendix F Integrating Linear RTRL Methods with PPO
When performing batch updates, as with PPO, the RTRL gradients used to update the recurrent parameters will be stale, as they were calculated during the interaction with the environment w.r.t old policy and value parameters. One solution to mitigate the gradient staleness is to go through the whole trajectory after each epoch update and re-compute the gradient traces. However, this can be computationally expensive. Algorithm 1, we provide the pseudocode for integrating RTRL methods with PPO with optional steps for re-running the network to update the RTRL gradient traces, the value targets, and the advantage estimates.
@@ Appendix F Integrating Linear RTRL Methods with PPO
In the next experiment, we investigate the effect of the gradient staleness in RTRL when combined with PPO and how this staleness interacts with the targets and advantage estimates. To understand this interaction, we evaluate all combinations of stale gradient and stale targets with increasing the number of epoch updates. We perform this analysis on the Ant-P environment from the Mujoco POMDP benchmark [30, 11, 26, 31]. Surprisingly, Figure 17 shows that using a stale gradient results in better performance with RTUs than re-computing the gradient traces. This performance improvement is also consistent when we increase the number of epochs from 4 to 8. It also shows that re-computing the value targets and advantage estimates has a minimal effect on the performance. We repeated the same experiments for NonLinear RTUs and Online LRU in Appendix F with consistent results.
One hypothesis for the superior performance of stale gradients is that the staleness is helping PPO maintain the trust region. We investigate this hypothesis by measuring the KL divergence between the policy used to collect the trajectory and the most recent policy. We use the following estimate for the KL divergence between the two policies as $(r-1)-\log(r)$, where $r=(\pi_{\theta_{new}})/(\pi_{\theta_{old}})$. The rightmost subplot of Figure 17 shows that at the beginning of learning, agents with stale gradients move away from the old policy more slowly than agents with fresh gradients; perhaps stale gradients helo maintaining the trust region.
@@
           Perform a gradient step with AutoDiff and correct it with the RTRL saved gradient as discussed in C.6.
@@
Appendix G Memory-Based Control Experiments
@@
In Figures 21 and 22, we set the truncation length for GRU and LRU to be $64$, which is larger than needed to solve the Mujoco POMDP tasks. These results show that even when the truncation length is larger than needed, RTUs still outperform T-BPTT baselines. We also show that the transformer-based models, GPT2, perform worse than RNN-based models. This is consistent with previous work suggesting that transformers might not be suitable for RL tasks [31].
Learning Rate Sensitivity:
@@
Figures 232425, and 26 show the learning rate sensitivity for all agents in the Mujoco POMDP benchmark.
Finally, we used the default hyper-parameters for PPO [39] indicated in Table 1 for all agents.
@@
Appendix H Compute resources

### v2

@@ 4.1 Ablation Study on Architectural Choices for RTUs and LRUs
In this first experiment, we investigate the impact of several architectural choices on the performance of RTUs and LRUs varying where nonlinearity is applied. We use a simple multi-step prediction task called Trace conditioning [39] inspired by experiments in animal learning. The agent’s objective is to predict a signal—called the Unconditional Stimulus (US)—conditioned on an earlier signal—the Conditional Stimulus (CS). The prediction is formulated as a return, $G_{t}\doteq\sum_{k=0}^{\infty}\gamma^{k}{\text{US}}_{t+k+1}$, where the agent’s goal is to estimate the value function for this return. More details on this environment and experimental settings are in Appendix F. Figure 1 summarizes the results.
@@ 4.2 Learning under resources constraints
In the following experiments, all agents consist of a recurrent layer followed by a linear layer generating the prediction. We measure performance of the agents online by calculating the Mean Square Return Errors (MSRE), which is the mean squared error between the agent’s prediction at time $t$ and $G_{t}$. In all the experiments, we used the Adam optimizer. We first ran each agent with different step sizes for $5$ runs and $2$ million steps. We then averaged the MSRE over the $2$ million steps and selected each agent’s best step size value. Finally, we ran the agents with the best step size value for $10$ runs, which we report here. We also report all agents’ step size sensitivity curves in Appendix F.3.
@@
We provide additional experiments comparing RTUs to two other approaches that use RTRL: online LRUs and a real-valued diagonal RNN in Appendix F.1.
@@ 5.1 Linear RTRL Methods in Incremental and Batch Settings
Integrating Linear RTRL Methods with PPO When performing batch updates, as with PPO, the RTRL gradients used to update the recurrent parameters will be stale, as they were calculated during the interaction with the environment w.r.t old policy and value parameters. One solution to mitigate the gradient staleness is to go through the whole trajectory after each epoch update and re-compute the gradient traces. However, this can be computationally expensive. In Appendix G, Algorithm 1, we provide the pseudocode for integrating RTRL methods with PPO with optional steps for re-running the network to update the RTRL gradient traces, the value targets, and the advantage estimates. We also performed an ablation study to investigate the effect of the gradient staleness in RTRL when combined with PPO, Appendix G. The results from the ablation study show that using a stale gradient results in better performance with RTUs and suggests that the staleness might help PPO maintain the trust region.
@@ 6 Experiments in Memory-Based Control
We evaluated our Linear and Non-linear RTUs against GRU, LRU, and Online LRU. All agents use PPO [40] as the control algorithm, and the difference between the agents is the recurrent component. For all agents, we fixed the number of parameters for the recurrent part to be $\sim 24$k. We tuned the learning rate for all agents in all environments and selected the best learning rate for each agent per environment. We also included a GPT2-transformer baseline. We followed the implementation details in previous work  [32], and used a GPT2 variant with $200$k parameters. We provide the results for GPT2 in Appendix H.
@@
When given the same number of parameters, RTU agents outperform other baselines in all environments in Figure 4. Furthermore, we show in Appendix H that even when increasing the truncation length of both GRU and LRU agents to use significantly longer history, they outperform RTUs in only one task. Of particular note is again that RTUs outperform online LRUs, highlighting again that our simple modifications have a large impact on performance in this online RL setting.
@@ 8 Acknowledgments
Appendix A Background on BackPropagation Through Time and Real-Time Recurrent Learning
A.1 BackPropagation Through Time
Applying the chain rule, we re-write Eq.5 as:
Once again, we need to consider the two pathways when evaluating $\frac{\partial\mathbf{h}_{t-1}}{\partial\boldsymbol{\psi}}$ in 8. For simplicity, let $\mathbf{J}_{t}\doteq\frac{\partial\mathbf{h}_{t}}{\partial\boldsymbol{\psi}}$, $\mathbf{B}_{t}=\frac{\partial\mathbf{f}(\mathbf{h}_{t-1},\mathbf{x}_{t},\boldsymbol{\psi})}{\partial\boldsymbol{\psi}}$, $\mathbf{C}_{t}=\frac{\partial\mathbf{f}(\mathbf{h}_{t-1},\mathbf{x}_{t},\boldsymbol{\psi})}{\partial\mathbf{h}_{t-1}}$, and re-write 8:
Writing $\frac{\partial\mathbf{h}_{t}}{\partial\boldsymbol{\psi}}$ using the results from 9:
A.1.1 Truncated-BackPropagation Through Time
Williams and Peng (1990) introduced Truncated-BackPropagation Through Time (T-BPTT) which solves the issue of increasing memory and computational complexities of BPTT. In T-BPTT, we specify a truncation length $T$, which controls the number of steps considered when calculating the gradient in 10. We now write the truncated version of 9 which takes into consideration the gradient from the last $T$ steps only:
Using results from 11, we then write the approximated gradient of the loss w.r.t the learnable parameters:
A.2 Real-Time Recurrent Learning
Williams and Zipser (1989) introduced the Real-time Recurrent Learning algorithm (RTRL) as a learning algorithm for continual recurrent learning. RTRL employs the recurrent formulation of the gradient in 8; instead of unrolling $\frac{\partial\mathbf{h}_{t-1}}{\partial\boldsymbol{\psi}}$ further back in time, RTRL saves its calculated value from the previous time step and use it later when needed. It is worth emphasizing that after the agent updates its parameters, the gradient information saved from previous time steps would be stale, i.e., calculated w.r.t old parameters, however, under the assumption of small learning rates, RTRL is known to converge. The gradient formulation of RTRL can be written as:
Appendix B More Details on Representability with Complex-valued Diagonal Recurrence
@@
B.1 Representability with Complex-valued Diagonals
@@
The above equivalence has only been used without any activation, namely in linear-recurrent units (LRUs) [34]. A natural question is if only the identity $f(\mathbf{x})=\mathbf{x}$ (and linear functions) satisfy this property of commuting with eigenvector matrices. Intuitively, this seems like the only option, as imagining a nonlinearity that commutes is hard. Surprisingly, for the more restricted case of symmetric $\mathbf{W}_{h}$, we can show a slightly more general class of activations can be used, proving an if-and-only-if relationship (see Appendix C). However, even for this restricted setting, this generalized class is limited and such activations unlikely to be preferable to a linear recurrence. We see this as a negative result, that suggests this equivalence only holds for the linear setting.
B.2 Complex Eigenvalues in Vanilla RNNs
We empirically investigate whether complex eigenvalues appear when training dense RNNs in a simple task. The goal is to show that the weight matrix, $\mathbf{W}_{h}$, is not a symmetric matrix, even in the simplest tasks. Hence, having only real-valued diagonals is too restrictive.
We used the Three State POMDP [42], depicted in Figure 7, for this experiment. In this task, the agent needs to remember one cue from the previous time-step ago, to make a prediction about the next time-step. The MDP has three states, $s_{1}$, $s_{2}$, and $s_{3}$, and no actions. If the agent is in either $s_{1}$ or $s_{2}$, it transitions to any of the three states with equal probability. However, if the agent is in $s_{3}$, it transitions to the state preceded by $s_{3}$. A sequence of observations would look like $1,3,1,2,2,3,2,\cdots$, and we ask the agent to predict the next observation.
Appendix C More on the Equivalence of Non-Linear RTUs and Dense RNNs
@@ D.2 Biased gradient when only using the real part of the hidden state
Appendix E Recurrent Trace Units
E.1 Empirical Analysis for different $r$ and $\theta$ parameterizations:
Since $r$ represents the magnitude of a complex number, then $r\in\mathbb{R}^{+}$, it is preferred to have $r\in(0,1]$ to avoid vanishing/exploding gradients as discussed in the previous section. Let $w_{r}$ be a learnable parameter which could directly represent $r$ or represent a function of $r$, we can enforce the constraints on $r$ in several ways:
  1. 1.
  2. 2.
  3. 3.
E.2 Real-Time Recurrent Learning for a Single Layer Linear RTUs
We can derive the following gradients:
where
We then get:
E.3 Single Layer Non-Linear RTUs Formulation:
We then get:
E.4 Complexity Analysis of RTUs
The saved gradient information has the following dimensions:
For the computational complexity, a forward pass according to 2 has a computational complexity of $\mathcal{O}(n+nd)$. Additionally, after doing the forward pass, the learner needs to update the saved gradient information according to equations 27 through 31 which has a computational complexity of $\mathcal{O}(n+nd)$. To summarize, using Real-Time Recurrent Learning with RTUs has linear computational and memory complexities.
E.5 Multi-Layers Recurrent Trace Units
where
E.6 Implementing RTRL within the reverse-mode automatic differentiation
Appendix F Additional Details on Trace Conditioning Experiments
@@
F.1 Additional Trace Conditioning Experiments
@@
Figure 11: Comparison to a block diagonal RNN.
The results in figure 11 indicate that these seemingly small differences between the diagonal RNNs can result in significantly different behavior. RTUs outperform online LRUs, with the differences discussed in-depth in Section 3.4. RTUs also outperform the block diagonal RNN. We emphasized using real-valued diagonals implicitly assumes symmetric matrices, but that is for a single real-value. This block diagonal has more representational capacity than the RTU. This result suggests it is beneficial for learning to enforce these constraints on the learnable parameters, that they correspond to the rotational representation of complex numbers.
On the role of RTRL in RTUs: To highlight the role of RTRL in RTUs, we evaluated the performance of both linear and non-linear RTUs with T-BPTT. In this experiment, we all agents use the same number of parameters; the only difference is whether they use RTRL or BPTT.
@@
F.2 Trace Conditioning Experiments Details
@@
F.3 Learning Rate Sensitivity
@@
Figure 15: RTUs used in the animal learning benchmark. The number in the label refers to the number of hidden units in the RTU.
@@
Appendix G Integrating Linear RTRL Methods with PPO
When performing batch updates, as with PPO, the RTRL gradients used to update the recurrent parameters will be stale, as they were calculated during the interaction with the environment w.r.t old policy and value parameters. One solution to mitigate the gradient staleness is to go through the whole trajectory after each epoch update and re-compute the gradient traces. However, this can be computationally expensive. In Algorithm 1, we provide the pseudocode for integrating RTRL methods with PPO with optional steps for re-running the network to update the RTRL gradient traces, the value targets, and the advantage estimates.
@@ Appendix G Integrating Linear RTRL Methods with PPO
In the next experiment, we investigate the effect of the gradient staleness in RTRL when combined with PPO and how this staleness interacts with the targets and advantage estimates. To understand this interaction, we evaluate all combinations of stale gradient and stale targets with increasing the number of epoch updates. We perform this analysis on the Ant-P environment from the Mujoco POMDP benchmark [31, 12, 27, 32]. Surprisingly, Figure 18 shows that using a stale gradient results in better performance with RTUs than re-computing the gradient traces. This performance improvement is also consistent when we increase the number of epochs from 4 to 8. It also shows that re-computing the value targets and advantage estimates has a minimal effect on the performance. We repeated the same experiments for NonLinear RTUs and Online LRU with consistent results in figures 19 and 20.
One hypothesis for the superior performance of stale gradients is that the staleness is helping PPO maintain the trust region. We investigate this hypothesis by measuring the KL divergence between the policy used to collect the trajectory and the most recent policy. We use the following estimate for the KL divergence between the two policies as $(r-1)-\log(r)$, where $r=(\pi_{\theta_{new}})/(\pi_{\theta_{old}})$. The rightmost subplot of Figure 18 shows that at the beginning of learning, agents with stale gradients move away from the old policy more slowly than agents with fresh gradients; perhaps stale gradients help with maintaining the trust region. However, this hypothesis still needs more investigation in future work.
@@
           Perform a gradient step with AutoDiff and correct it with the RTRL saved gradient as discussed in E.6.
@@
Appendix H More Details on the Memory-Based Control Experiments
@@
In Figures 22 and 23, we set the truncation length for GRU and LRU to be $64$, which is larger than needed to solve the Mujoco POMDP tasks. These results show that even when the truncation length is larger than needed, RTUs still outperform T-BPTT baselines. We also show that the transformer-based models, GPT2, perform worse than RNN-based models. This is consistent with previous work suggesting that transformers might not be suitable for RL tasks [32].
Learning Rate Sensitivity: Figures 242526, and 27 show the learning rate sensitivity for all agents in the Mujoco POMDP benchmark. Finally, we used the default hyper-parameters for PPO [40] indicated in Table 1 for all agents.
@@
Appendix I Compute resources